### v2

@@ 3.2 Conditional denoising in map generation
Map generation models rely solely on remote sensing imagery tiles as inputs to produce corresponding spatial scene map tiles.This process can be mathematically expressed as follows:
@@
where $con$ and $rsm$ represent scale information and cascading reference conditions. Both serve as pivotal components for encoding map scale information. In the context of this methodology, the terms are defined as follows:
@@
  • Scale Information: The cross-modal constraints of scale features correspond to the map’s level. These encapsulate the spatial resolution, scale, and textual representation of geographic objects associated with the specified level of the generated map.
  • Cascade Reference: The visual representation of map sheets and geographic object information from lower levels, acts as cascading feature constraints.
During the training phase, any useful prior knowledge can be utilized to provide content guidance. In the testing phase, multi-scale reconstruction results are derived using scale information and cascading reference. By decomposing cartographic reference information into scale and cascading features, this modeling approach enables improved reconstruction guidance, delivering higher-quality generation results with enhanced scale representation and stronger semantic fidelity.
@@
Figure 3: The SGDM architecture leverages a Variational Autoencoder (VAE) to transition the diffusion and reverse processes from pixel space to latent space. During training, the latent representation of the target map, $z_{0}$, undergoes a progressive transformation into $z_{t}$ through the diffusion process, followed by denoising achieved via a U-Net network. Two bespoke modules are introduced to further refine the denoising process: the Map Feature Encoder (MFEncoder) and the Scale Feature Adapter (SFAdapter). The MFEncoder integrates information derived from remote sensing imagery and cascading references to construct the conditional feature, $F_{\mathrm{cond}}$. Expanding on $F_{\mathrm{cond}}$, the SFAdapter produces multi-scale features that are subsequently merged with the outputs of corresponding U-Net layers through element-wise addition.
3.3 Scale-guided diffusion model
In order to address the challenges associated with multi-scale map generation from remote sensing images in real-world scenarios, we propose a novel scale-unified map generation model called the Scale-guided Diffusion Model (SGDM). Unlike previous single-scale map generation models, SGDM can simultaneously process inputs from remote sensing image tiles, target map tiles, and cascading reference tiles. This collaborative approach enables the generation of map tiles customized to specific scales or levels. The core architecture of SGDM is illustrated in Fig. 3. The SGDM framework is composed of four integral components: the pre-trained Variational Auto-Encoder (VAE), the denoising U-shaped network (U-Net) with skip connections, the Map Feature Encoder (MFEncoder), and the Scale Feature Adapter (SFAdapter). The VAE encodes and decodes high-resolution images, thereby transitioning the conventional DDPM diffusion and denoising process from pixel space to latent space, enhancing training stability and reducing computational overhead. Concurrently, the U-Net, functioning as a conditioned noise predictor, iteratively predicts and removes noise to achieve clean data. It is constructed with multiple encoding, middle, and decoding layers, most of which are composed of residual blocks, cross-attention, and self-attention modules. The MFEncoder integrates prior information from the style guide map, remote sensing image, and scale reference map to produce a conditional feature map enriched with semantic and stylistic details. Finally, the Scale Feature Adapter (SFAdapter) converts these conditional features into multi-scale features (about the image domain scale rather than the geographic domain scale) that are compatible with the U-Net architecture. These features are incorporated into the outputs of the U-Net’s encoding and decoding layers on an element-wise basis, ensuring seamless feature integration and effective guidance throughout the generation process.
3.3.1 Scale information guidance
@@
At each generation stage, the previously generated lower-level tiles, along with time step and scale information embeddings are utilized as conditional input variables. For the lower-level tile input — cascading reference map $x_{0}^{(k)}$ — its dimensions do not align with $x_{t}^{(k+1)}$. To address this, and inspired by [29] and [69], we redesigned the network architecture. First, as illustrated in Fig. 3, the encoder $E_{lr}$ is applied to encode the lower-level map tiles $x_{0}^{(k)}$. Then, a series of upsampling and convolutional layers $f_{up}$ are employed to align the feature map dimensions with $x_{t}^{(k+1)}$. Finally, the features are fused by concatenating $x_{t}^{(k+1)}$ and the feature maps along the channel dimension:
@@
For the scale information $s$ in text modality data, we perform cross-modal encoding using pre-trained CLIP model [70, 71], with the CLIP scale encoder denoted as $f_{\theta}$ (Fig. 3). This approach allows us to obtain the embedding vector of scale information $e_{s}^{(k)}=f_{\theta}(s)$ at each generation stage, capturing the cross-modal representation of scale information corresponding to the spatial context of the generated map. This embedding subsequently guides the map generation process through the cross-attention mechanism [72] during denoising:
@@
Here, $\theta_{t}$ is a learnable parameter.
@@
Based on (7), while keeping the forward process unchanged, we construct a conditional denoising model:
@@
Through this design, we can ultimately generate map tiles of varying scales in a self-cascading manner, encoded with information on resolution, temporal steps, and scale.
Figure 4: Detailed structure of the proposed Multi-branch Mapping Feature Encoder (MFEncoder): it can generate the conditioned features $F_{\mathrm{cond}}$ that contain cartographic semantic information by utilizing the content information of the reference map at multiple scales through the SPADE module.
3.3.2 Map feature encoder
To facilitate the successful amalgamation of cascading information from cascading reference tiles into remote sensing image tiles, we have developed a multi-branched map feature encoder (MFEncoder), illustrated in Fig. 3. This encoder integrates semantic information from cascading reference tiles into remote sensing image tiles across various scales. As illustrated in Fig. 4(a) convolutional and down-sampling blocks are employed in one branch to extract multi-scale features from the cascading reference tiles. The multi-scale features are derived from the remote sensing image tiles through the Basic Block in another branch. To achieve the fusion of these features, we have incorporated the SPADE (Spatially-Adaptive Normalization) module at each scale. In our design, the input to the SPADE block comprises feature maps from remote sensing image tiles, abundant in coarse texture and structural information. As illustrated in Fig. 4(b), our enhanced SPADE module distinguishes itself from the original design by utilizing features from both remote sensing and cascading reference tiles to compute the scaling factor $\gamma$ and the offset term $\beta$. This approach enables a flexible and high-fidelity integration of content-guided information into the features of remote sensing images. The entire process is detailed as follows:
$\gamma^{i}=\mathrm{Conv}_{\gamma}\left(\mathrm{Conv_{share}}\left(\left[F_{\mathrm{rs}}^{i},F_{\mathrm{cr}}^{i}\right]\right)\right)$ (16)
$\beta^{i}=\mathrm{Conv}_{\beta}\left(\mathrm{Conv_{share}}\left(\left[F_{\mathrm{rs}}^{i},F_{\mathrm{cr}}^{i}\right]\right)\right)$ (17)
$F_{\mathrm{rs}}^{i+1}=\gamma^{i}\frac{F_{\mathrm{rs}}^{i}-\mu^{i}}{\sigma^{i}}+\beta^{i}$ (18)
In this context, $F_{\mathrm{rs}}^{i}$ and $F_{\mathrm{cr}}^{i}$ denote the features of terrestrial objects in remote sensing imagery and the features of the cascading reference map at the $i$-th layer, respectively. $\mathrm{Conv_{share}}$ represents the shared convolutional module, while $\mathrm{Conv}{\gamma}$ and $\mathrm{Conv}{\beta}$ are the convolutional layers used to predict the scaling factor $\gamma^{i}$ and the bias term $\beta^{i}$. Here, $\gamma^{i}$ and $\beta^{i}$ indicate the spatially varying scaling factor and bias term at the $i$-th layer, whereas $\mu^{i}$ and $\sigma^{i}$ correspond to the mean and standard deviation of all input features within the mini-batch. By stacking content-guiding modules, the features of the cascading reference map are progressively integrated into the terrestrial object features of remote sensing imagery, enabling more precise map reconstruction at a semantic level. Figure 4(c) illustrates the network architecture of the basic block, convolutional block, and downsampling block.
3.3.3 Scale feature adapter
The MFEncoder excels in integrating texture and structural features from remote sensing image tiles, alongside content features from cascading reference tiles, ultimately producing the conditional feature $F_{\mathrm{cond}}$. Given that $F_{\mathrm{cond}}$ encompasses the majority of external information requisite for map generation, the function of the Scale Feature Adapter (SFAdapter), illustrated in Fig. 3, is primarily to generate multi-scale features and align them with the pre-existing knowledge of the pre-trained stable diffusion model.
Consequently, the architecture of the SFAdapter module is inherently straightforward, primarily comprising stacked convolutional layers, residual blocks, and down-sampling blocks (Fig. 4(c)). The SFAdapter module progressively increases the channel count of the feature maps while simultaneously diminishing their spatial dimensions, resulting in three distinct sets of conditional features. These features are then integrated into the output of the U-Net network layer by layer, on an element-wise basis, to facilitate the generation process of the stable diffusion model.
Figure 5: Examples of RS-Map tile sample pairs from levels 14 to 18 in the CSCMG dataset are provided. The dataset covers a broad range of cross-scale scenarios, including forests, rivers, streets, residential areas, industrial areas, bodies of water, plazas, and more.
@@ 3.4 Loss function
As illustrated in Fig. 3, the trainable parameters of the Scale-Guided Diffusion Model (SGDM) are determined by the following loss function:
@@
This equation, $rs$ and $cr$, represents the input remote sensing image tiles and cascading reference tiles. Here, $\epsilon$ denotes randomly sampled Gaussian noise, while $t$ signifies the integer timestamp sampled from the interval $[0,T]$. In this context, $\epsilon_{\theta}$ refers to the denoising U-Net model.
@@ 3.5 Evaluation metrics
We adhere to established practices in generative cartography [17, 13] by employing image quality evaluation metrics to comprehensively assess the performance of map generation methods. The realism of the generated maps is quantified using the Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM), calculated on the Y channel within the YCbCr color space [73]. Furthermore, the Fréchet Inception Distance (FID) is utilized to evaluate the distributional divergence between real and generated maps [74].
While semantic segmentation metrics (e.g., mIoU) could offer deeper insights, our CSCMG dataset focuses on large-scale, unpaired image-to-image translation without pixel-level labels. To address the limitations of pixel or distribution-based metrics in assessing semantic-level cartographic features, we employ the Map Feature Perception Metric (MFP). This metric complements traditional evaluations by explicitly measuring global semantic alignment and spatial consistency between synthesized and reference maps. The MFP leverages a self-supervised Vision Transformer (ViT) [75] pretrained with DINO [76] to extract deep features, comprising two components:
  • Global Feature Evaluation: Captures holistic cartographic attributes (e.g., scene categories, element distributions) using the [CLS] token embeddings from the final ViT layer:
    $\mathcal{L}_{G}=1-\text{MSE}(\text{CLS}_{o},\text{CLS}_{t}),$ (20)
@@
  • Spatial Similarity Evaluation: Quantifies local geometric relationships via self-attention keys from the final ViT layer:
    $\mathcal{L}_{S}=\left\|S(I_{o})-S(I_{t})\right\|_{F},$ (21)
@@
$\text{MFP}=\lambda_{1}\mathcal{L}_{G}+\lambda_{2}\mathcal{L}_{S},$ (22)
with $\lambda_{1}=10$ and $\lambda_{2}=1$ empirically determined. Lower MFP values indicate superior semantic fidelity and spatial coherence.
We combine FID, PSNR, and SSIM with MFP to evaluate structural coherence, style fidelity, and semantic plausibility holistically.
@@ 4.1.1 Datasets
In this pursuit, we have established a novel benchmark dataset CSCMG to validate the proposed C2GM framework and stimulate future research in large-scale remote sensing image super-resolution. The data were meticulously collected from the Glasgow area in the UK, ensuring that the sampling regions for the test and training tile sets were distinctly delineated and non-overlapping. Representative image pairs are illustrated in Fig. 5.
Table 1: Detailed information for the CSCMG dataset.
Level Tile pair number Scale Resolution (m/pixel) Example features
@@
14 4870 280 1:35000 9.555 village, or suburb
15 15106 292 1:15000 4.777 small road
16 30040 300 1:8000 2.389 street
17 40555 300 1:4000 1.194 block, park, addresses
18 45000 300 1:2000 0.597 some buildings, trees
@@
Figure 6: Examples of cascade references from levels 16 to 18 in the CSCMG dataset. The dataset provides 2X or 4X cascade references, with image resolutions of 128 and 64.
CSCMG comprises pairs of remote sensing and map tiles across various scales, ranging from levels 14 to 18, as detailed in Table 1. The dataset includes 137,042 sample tiles, with 135,571 allocated for training and 1,471 designated for testing. All tiles were sourced from Google Maps’ free tile service. The extensive multi-scale tile samples within the dataset enable robust validation of multi-scale map generation capabilities.
Tiles span from level 18 down to level 14, encompassing the entire spatial extent of all sub-level tiles structured through a quadtree division. This design enhances both the training and testing processes of the cascade map generation model. The dataset offers cascade training data at 2X and 4X, providing cascading references covering either one or two levels. As illustrated in Fig. 6, the image resolutions of the reference tiles are either 128 or 64, upscaled to 256 to ensure uniform tile sizes while preserving the spatial resolution of the original references. This approach facilitates the validation of the efficacy of multi-scale cascading generation. Bi-cubic resampling is employed to maintain a 2X or 4X spatial resolution difference between the map tiles and cascading reference tiles.
To evaluate our method’s multi-scale map generation capabilities, we utilize the multi-scale map generation dataset MLMG constructed by the state-of-the-art (SOTA) method LACG [17]. The MLMG dataset assesses multi-scale map generation techniques and includes tiles from diverse urban environments such as forests, rivers, streets, residential areas, industrial zones, water bodies, and squares. The dataset comprises samples collected from cities in the United States and China, reflecting significant differences in their respective images and maps. In China, roads are typically wider with lower distribution density, lacking a clear parallel-vertical relationship, whereas in the United States, roads are generally straight and densely packed, with the ground neatly segmented into blocks. The dataset from the U.S. city is MLMG-US, while the dataset from the Chinese city is MLMG-CN. Both MLMG-US and MLMG-CN contain equal samples and levels, featuring multi-scale remote sensing images paired with corresponding map tiles and element labels.
@@ 4.1.2 Implementation details
Our proposed C2GM framework is based on the Stable Diffusion 2.1-base model. The SGDM model is implemented using the PyTorch framework, and all experiments are conducted on two NVIDIA GeForce RTX 3090 GPUs. The training process leverages cascaded map tiles spanning levels 14 to 18 from the CSCMG dataset.
The training procedure is divided into two distinct stages. In the first stage, the Stable Diffusion model is fine-tuned using the loss function defined in Section 3.4. During this phase, the parameters of the Map Feature Encoder (MFEncoder), Scale Feature Adapter (SFAdapter), and the linear layers of the U-Net are optimized. The model consists of 152 million trainable parameters, with a total size of 1.4 billion, and employs a hierarchical architecture that reduces computational redundancy through parameter sharing across scales. Training proceeds over 200,000 steps with a batch size of 20, utilizing the AdamW optimizer with a learning rate of $5\times 10^{-5}$, and achieves convergence within 480 GPU-hours using dual RTX 3090 GPUs.
C2GM demonstrates enhanced computational efficiency through latent space compression and cascaded generation. A single 256×256 tile is generated in 1.2 seconds on two RTX 3090 GPUs, achieving a total of 3.2 trillion floating-point operations per iteration, a 42% reduction compared to standard diffusion models. For practical emergency mapping at a 1:50,000 scale, covering zoom levels 10 to 16 and encompassing 441 tiles, the end-to-end generation completes in approximately 9 minutes. Although diffusion models inherently exhibit higher complexity than GANs due to iterative denoising, our architectural optimizations, such as multi-scale feature reuse and CLIP-guided early stopping, effectively mitigate computational overhead.
Figure 7: Visualized results of map generation on the MLMG-US test set using our method and analogous approaches are presented. However, due to the limitation of reference information, only the results for levels 17 and 18 are available in section (k) of the figure. Our results demonstrate superior accuracy in map scale representation and enhanced visual continuity of the combined tile map.

### v3

@@ 3.2 Conditional denoising map generation
Map generation relies solely on remote sensing imagery tiles as inputs to produce corresponding spatial scene map tiles. This conditional denoising process can be mathematically expressed as follows:
@@
where $con$ and $rsm$ represent scale information and cascading reference conditions. Both serve as pivotal components for encoding cartographic features. In the context of this methodology, the terms are defined as follows:
@@
  • Scale Information: The spatial resolution, map level, and feature types associated with the specified scale of the generated map.
  • Cascade Reference: The smaller-scale map tiles produced in the previous stage serve as geographic feature cascading constraints.
@@
Figure 3: The SCGM architecture leverages a VAE to transition the diffusion and reverse processes from pixel space to latent space. During training, the latent representation of the target map, $z_{0}$, undergoes a progressive transformation into $z_{t}$ through the diffusion process, followed by denoising achieved via a U-Net network. Two bespoke modules are introduced to further refine the denoising process: MFEncoder and SFAdapter. The MFEncoder integrates information derived from remote sensing imagery and cascading references to construct the conditional feature, $F_{\mathrm{cond}}$. Expanding on $F_{\mathrm{cond}}$, the SFAdapter produces multi-scale features that are subsequently merged with the outputs of corresponding U-Net layers through element-wise addition.
3.3 Scale information guidance
At each generation stage, previously generated smaller-scale tiles, along with scale information embeddings, serve as conditional inputs. Since the cascading reference map $x_{0}^{(k)}$ differs in resolution from the current input $x_{t}^{(k+1)}$, we adapt the architecture following principles from (Saharia et al., 2022c; Li et al., 2022). As shown in Figure 3, the encoder $E_{lr}$ first extracts features from $x_{0}^{(k)}$, which are then upsampled and transformed via a module $f_{up}$ comprising convolutional and interpolation layers to match the spatial dimensions of $x_{t}^{(k+1)}$. The resulting features are concatenated with $x_{t}^{(k+1)}$ along the channel dimension for fusion.
@@ 3.3 Scale information guidance
At each generation stage, the previously generated small-scale tiles, along with time step and scale information embeddings are utilized as conditional input variables. For the smaller-scale tile input — cascading reference map $x_{0}^{(k)}$ — its dimensions do not align with $x_{t}^{(k+1)}$. To address this, and inspired by (Saharia et al., 2022c) and (Li et al., 2022), we redesigned the network architecture. First, as illustrated in Figure 2, the encoder $E_{lr}$ is applied to encode the smaller-scale map tiles $x_{0}^{(k)}$. Then, a series of upsampling and convolutional layers $f_{up}$ are employed to align the feature map dimensions with $x_{t}^{(k+1)}$. Finally, the features are fused by concatenating $x_{t}^{(k+1)}$ and the feature maps along the channel dimension:
@@
For the scale information $s$ in text modality data, we perform cross-modal encoding using pre-trained CLIP model (Radford et al., 2021; Cherti et al., 2023), with the ScaleEncoder denoted as $f_{\theta}$ (Figure 3). This approach allows us to obtain the embedding vector of scale information $e_{s}^{(k)}=f_{\theta}(s)$ at each generation stage, capturing the cross-modal representation of scale information corresponding to the spatial context of the generated map. This embedding subsequently guides the map generation process through the cross-attention mechanism (Vaswani et al., 2017) during denoising:
@@
where $\theta_{t}$ is a learnable parameter.
@@
Based on Equation (3), while keeping the forward process unchanged, we construct a conditional denoising model:
@@
Through this design, we can ultimately generate map tiles of varying scales in a self-cascading manner, encoded with information on temporal steps, and scale information.
Figure 4: Detailed structure of the MFEncoder, SPADE module and basic blocks.
3.4 Map feature encoder
To enable effective integration of semantic cues from cascading reference tiles into remote sensing image tiles, we propose a multi-branch map feature encoder MFEncoder, as shown in Figure 3. This encoder incorporates multi-scale semantic information from the reference tiles into the remote sensing features. As illustrated in Figure 4(a), one branch extracts hierarchical features from the reference tiles using convolutional and downsampling blocks, while the other branch processes remote sensing image tiles through a series of basic blocks.
We employ the SPADE block (Park et al., 2019) at each stage to facilitate feature fusion across scales. The SPADE block utilizes remote sensing features—characterized by coarse textures and structural details—as its primary input in our design. As illustrated in Figure 4(b), features from both remote sensing and reference tiles are jointly leveraged to compute the scale ($\gamma$) and bias ($\beta$) parameters. This design enables more adaptive and semantically aligned fusion, enhancing the fidelity of content-aware feature integration (Wang and Sun, 2025). The structure of each module, encompassing basic, convolutional, and downsampling blocks, is depicted in Figure 4(c).
3.5 Scale feature adapter
The SFAdapter module (Figure 3) transforms the dimensions of scale-mapped features received from the encoder to generate scale-specific feature representations compatible with the latent space organization of the pre-trained diffusion model. It synthesizes the final comprehensive conditional features $F_{\mathrm{cond}}$ by integrating texture and structural patterns extracted from image patches by the MFEncoder with contextual attributes obtained from the cascading reference blocks. The SFAdapter demonstrates exceptional capability in fusing multimodal features.
Specifically, the architecture of the SFAdapter module is straightforward, as detailed in Figure 4(c), employing sequential processing stages through convolutional operations, residual connections, and progressive downsampling. The module systematically expands feature channel capacity through successive transformations while contracting spatial resolution, producing three hierarchical conditional features at varying scales. These multi-scale representations are subsequently merged with corresponding hierarchical features in the U-Net decoder through channel-wise concatenation and additive fusion, thereby enabling spatial-adaptive modulation of the diffusion model’s generation pathway.
@@ 3.6 Loss function
The trainable parameters of the SCGM are determined by the following loss function:
@@
this eqnarray, $rs$ and $cr$ represent the input remote sensing image tiles and cascading reference tiles. Here, $\epsilon$ denotes randomly sampled Gaussian noise, while $t$ signifies the integer timestamp sampled from the interval $[0,T]$. In this context, $\epsilon_{\theta}$ refers to the denoising U-Net model.
Figure 5: Examples of RS-Map tile sample pairs from scales 1:35,000 to 1:2,000 in the CSCMG dataset are provided, covers a broad range of cross-scale scenarios.
@@ 3.7 Evaluation metrics
We adhere to established practices in generative mapping (Fu et al., 2023, 2022) by employing quality evaluation metrics to comprehensively assess the performance of map generation methods. The realism of the generated maps is quantified using the PSNR and SSIM, calculated on the Y channel within the YCbCr color space (Wang et al., 2004). Furthermore, the FID is utilized to evaluate the distributional divergence between real and generated maps (Heusel et al., 2017).
While semantic segmentation metrics (e.g., mIoU) could offer deeper insights, our method focuses on large-scale, paired image-to-image translation without pixel-level semantic labels. To address the limitations of pixel or distribution-based metrics in assessing semantic-level cartographic features, we employ the map feature perception metric MFP (Sun and Bai, 2025). This metric complements traditional evaluations by explicitly measuring global semantic alignment and spatial consistency between synthesized and reference maps. The MFP leverages a self-supervised Vision Transformer (ViT) pretrained with DINO (Caron et al., 2021) to extract deep features, comprising two components:
Global feature evaluation captures holistic cartographic attributes (e.g., scene categories, element distributions) using the [CLS] token embeddings from the final ViT layer:
$\mathcal{L}_{G}=1-\text{MSE}(\text{CLS}_{o},\text{CLS}_{t})$ (13)
@@
Spatial similarity evaluation quantifies local geometric relationships via self-attention keys from the final ViT layer:
$\mathcal{L}_{S}=\left\|S(I_{o})-S(I_{t})\right\|_{F}$ (14)
@@
$\text{MFP}=\lambda_{1}\mathcal{L}_{G}+\lambda_{2}\mathcal{L}_{S}$ (15)
with $\lambda_{1}=10$ and $\lambda_{2}=1$ empirically determined. Lower MFP values indicate superior semantic fidelity and spatial coherence. A detailed explanation of the interpretability and spatial consistency of the MFP metric is provided in (Sun and Bai, 2025).
@@ 4.1.1 Datasets
To address limitations in existing multi-scale map datasets, we present the cross-scale cascade map generation dataset CSCMG, comprising real-world remote sensing images and tile maps from Glasgow and London, UK. CSCMG pairs multi-scale tile maps with corresponding imagery, sampled from non-overlapping Glasgow regions for training/testing and continuous London areas for seamless generation evaluation. Each pair includes scale information (map scale, level, resolution), enabling cross-scale validation. Representative examples are shown in Figure 5 and the detailed information shown in Table 1.
Table 1: The detailed information of the CSCMG dataset, including the quantity of RS-Map tile sample pairs across various scales, as well as scale information (map scale, level, resolution, and feature types).
Level Tile pair number Scale Resolution (m/pixel) Feature types
@@
14 4870 280 1:35,000 9.555 village, or suburb
15 15106 292 1:15,000 4.777 small road
16 30040 300 1:8,000 2.389 street
17 40555 300 1:4,000 1.194 block, park, addresses
18 45000 300 1:2,000 0.597 some buildings, trees
@@
Figure 6: Examples of cascade references from scales of 1:8,000 to 1:2,000 in the CSCMG dataset. The dataset provides 2X or 4X cascade references, with image resolutions of 128 and 64.
The CSCMG dataset comprises 137,042 paired remote sensing and map tiles spanning scales from scales 1:35,000 to 1:2,000, sourced from Google Maps’ free tile service. These tiles are organized through quadtree division to hierarchically span spatial extents ranging from scales of 1:35,000 (the coarsest) to 1:2,000 (the finest), including 135,571 tile pairs for training and 1,471 for testing. As illustrated in Figure 6, cascade references at 2X and 4X scales provide multi-scale constraints: reference tiles (64 or 128 resolution) are upscaled to 256×256 via bi-cubic resampling, ensuring uniform tile sizes while preserving spatial resolution relationships. This design verifies multi-scale cascading generation efficacy, particularly in hierarchical feature alignment and seamless map tile synthesis.
To evaluate our method’s multi-scale map generation capabilities, we utilize the multi-scale map generation dataset MLMG constructed by the SOTA method LACG (Fu et al., 2023). The MLMG comprises samples collected from cities in the US and China, reflecting significant differences in their respective images and maps. In China, roads are typically wider with lower distribution density, lacking a clear parallel-vertical relationship, whereas in the United States, roads are generally straight and densely packed, with the ground neatly segmented into blocks.
@@ 4.1.2 Implementation details
The proposed SCGM framework is developed on the Stable Diffusion pretrained model (SD-2.1-base) and implemented using PyTorch. All experiments are performed on dual NVIDIA GeForce RTX 3090 GPUs, utilizing cascaded map tiles and remote sensing imagery from scales 1:35,000 to 1:2,000 of the CSCMG dataset.
In the training process, the pretrained model is fine-tuned using the loss function described in Section 3.6. Optimization targets include the Map Feature Encoder (MFEncoder), Scale Feature Adapter (SFAdapter), and the linear layers of the U-Net. The model contains 152 million trainable parameters out of a total of 1.4 billion and adopts a hierarchical structure to reduce redundancy via cross-scale parameter sharing. Training is conducted for 200,000 steps with a batch size of 20, using the AdamW optimizer with a learning rate of $5\times 10^{-5}$, and converges within 480 GPU-hours on two RTX 3090 GPUs.
Figure 7: Generated results on the MLMG-US test set using SCGM and baseline methods, highlighting SCGM’s superior seamless tile synthesis and geographic detail preservation across scales.